# TROVE: A Challenge for Fine-Grained Text Provenance via Source Sentence Tracing and Relationship Classification

### v1

@@ Abstract
LLMs have achieved remarkable fluency and coherence in text generation, yet their widespread adoption has raised concerns about content reliability and accountability. In high-stakes domains such as healthcare, law, and news, it is crucial to understand where and how the content is created. To address this, we introduce the Text pROVEnance (TROVE) challenge, designed to trace each sentence of a target text back to specific source sentences within potentially lengthy or multi-document inputs. Beyond identifying sources, TROVE annotates the fine-grained relationships (quotation, compression, inference, and others), providing a deep understanding of how each target sentence is formed. To benchmark TROVE, we construct our dataset by leveraging three public datasets covering 11 diverse scenarios (e.g., QA and summarization) in English and Chinese, spanning source texts of varying lengths (0–5k, 5–10k, 10k+), emphasizing the multi-document and long-document settings essential for provenance. To ensure high-quality data, we employ a three-stage annotation process: sentence retrieval, GPT provenance, and human provenance. We evaluate 11 LLMs under direct prompting and retrieval-augmented paradigms, revealing that retrieval is essential for robust performance, larger models perform better in complex relationship classification, and closed-source models often lead, yet open-source models show significant promise, particularly with retrieval augmentation.
@@ 1 Introduction
Specifically, we first select examples with multi-sentence outputs to ensure sufficient sources and categorize data by original scenarios (question answering or summarization), languages (Chinese or English), input text length (0-5k, 5k-10k, or 10k+), and number of input documents (single-document or multi-document), aiming for a balanced distribution across each dimension. Next, we employ multiple information retrieval methods to recall the candidate source sentences for each sentence in the target text. Then, we use crafted prompts to guide GPT-4 in preliminary annotation, aiming to identify the sources of target sentences from retrieved candidate sentences and classify the target-source relationships into quotation (verbatim or partial copy), compression (summarization or paraphrase), inference, and others (e.g., negation). Each target sentence can be traced to multiple source sentences, and different relationships may apply simultaneously. For instance, a target sentence sourced from [a, b, c] might have [a, b] labeled as compression and [c] as quotation. Finally, we conduct a human review, requiring annotators to verify each annotation by considering both the source document(s) and the results produced by GPT-4.
@@ 3 Task Formulation
Figure 2: Overview of our data annotation. (a) Sentence Retrieval: selecting candidate provenance sentences using multiple retrievers; (b) GPT-4 Provenance: automatically annotating provenance relationships based on retrieved sentences; (c) Human Provenance: reviewing and refining GPT-4’s annotations while independently checking source documents to identify missing provenance sentences. Di-j denotes the j-th sentence in the i-th document.
@@ 4.2 Data Annotation
We employ GPT-4 to alleviate the manual annotation workload, as shown in Figure 2. For each target sentence $t_{i}$, the annotation procedure consists of three steps: (a) sentence retrieval, (b) GPT-4 provenance, and (c) human provenance.
Sentence Retrieval. Due to the lengthy input text, GPT-4 may ignore key sentences when directly tracing provenance through long passages. To mitigate this, we first retrieve candidate provenance sentences using the target sentences as queries and then perform provenance based on these sentences. We have presented ablation studies to validate this approach. Moreover, different retrievers capture diverse semantic features. To maximize the recall rate of candidate provenance sentences, we aggregate the results from $M$ distinct retrievers. Each retriever selects the top-$k$ most relevant sentences, denoted as $R_{i}(D,t_{i})$, forming the following set of candidate provenance sentences:
@@
GPT-4 Provenance. Based on the candidate provenance sentences, GPT-4 conducts fine-grained annotation and classifies the provenance relationship types, as depicted in Figure 2(b). The detailed prompt is provided in the appendix.
Human Provenance. Annotators review GPT-4’s results to verify the provenance sentences and their corresponding relationship types. Noting that GPT-4 can ignore critical details, the annotators examine the document collections to address any omissions. As illustrated in Figure 2(c), the sentences “D2-2” and “D2-4”, which contain “12,000 books” and “digital reading machine” respectively, exhibit a strong connection to the target sentence, yet GPT-4 fails to identify them. Therefore, the annotators will incorporate these two sentences into the final analysis. We invite 8 graduate students to spend about 510 hours annotating the provenance of 4,388 English and 811 Chinese sentences, costing an average of \$0.20 per sentence.
@@ 4.4 Consistency Analysis
#Docs Lang Trace Type GPT-4
@@
To ensure dataset quality, 10% of the examples are assigned to different annotators for consistency assessment. We evaluate annotator agreement from three perspectives: (1) tracing provenance sentences, (2) classification of relationship types, and (3) determination of necessary corrections to GPT-4’s provenance sentences. To quantify agreement, we use Fleiss’ Kappa (Falotico and Quatto, 2015) to measure the reliability across multiple annotators. The results, presented in Table 2, demonstrate that the annotation process is reliable.
@@ 5.1 Experimental Setup
Besides, during GPT provenance, we find that an initial retrieval step significantly enhances GPT’s recall rate. Thus, each model is evaluated under two approaches: (1) direct prompting tracing, where the model processes the input directly, and (2) retrieval-augmented tracing, where retrieval is performed first, followed by tracing based on the retrieved results.
@@ 5.3 Evaluation Results
Vicuna-7b DP 06.80 23.22 10.50 02.37 08.53 10.61 06.78 23.10 10.44 02.38 08.43 10.52 10.51
RA 27.14 41.76 32.50 9.79 17.22 34.20 29.10 40.78 33.64 10.39 15.76 34.74 33.77
LLama3-8b DP 05.16 16.63 06.97 02.03 07.21 07.21 05.45 15.50 06.43 01.84 06.20 06.81 06.86
RA 43.74 38.19 40.61 22.49 19.51 40.32 49.81 35.04 41.07 25.42 18.40 41.74 40.94
Chatglm-6b DP 00.02 00.04 00.02 00.01 00.00 00.01 00.04 00.02 00.03 00.01 00.01 00.04 0.02
RA 03.68 04.06 03.84 01.50 01.76 03.96 11.97 03.98 05.93 04.85 01.73 06.12 4.96
Qwen1.5-7b DP 06.47 36.63 10.67 01.26 08.69 11.01 06.23 36.36 10.33 01.17 07.66 10.56 10.64
RA 37.22 52.87 43.05 09.31 16.57 47.82 35.68 52.02 41.62 08.85 14.53 44.32 44.20
Qwen2.5-7b DP 08.88 49.56 14.77 02.89 15.49 14.66 07.67 49.50 12.94 01.93 13.56 13.08 13.86
RA 41.96 71.92 52.23 14.73 28.56 55.06 39.50 69.72 49.65 12.32 24.55 52.22 52.29
Qwen1.5-14b DP 12.50 38.97 16.84 3.71 13.35 17.32 13.36 37.51 16.93 03.65 11.32 17.12 17.05
RA 45.07 53.68 48.33 14.06 20.96 53.47 47.99 51.65 49.20 14.50 17.90 52.27 50.82
Qwen2.5-14b DP 29.77 52.25 36.12 15.11 27.37 36.22 29.50 49.55 33.01 14.37 24.74 33.00 34.59
RA 54.60 50.24 51.99 29.22 27.43 52.43 64.68 47.02 54.23 33.49 24.80 54.82 53.37
DeepSeek-V3 (671B) DP 44.79 56.56 49.80 21.88 28.40 50.02 39.54 54.31 44.95 17.46 25.61 45.59 47.59
RA 49.17 55.85 51.94 26.10 31.20 52.96 50.75 53.76 51.93 26.17 28.54 52.74 52.39
@@
Impact of Retrieval-Augmented Tracing vs. Direct Prompting Tracing. Across almost all models, retrieval-augmented tracing outperforms direct prompting in F1 scores, often by a large margin. For example, Qwen2.5-14B’s F1 jumps from 34.59 to 53.37 with retrieval, while ChatGLM-6B, which nearly fails in direct prompting tracing with an F1 of 0.02, improves to 4.96. Even closed-source models show the same trend, as Gemini-1.5-Pro significantly improves from an F1 of 12.29 to 63.36 with retrieval. It suggests that retrieval helps overcome context-length limits and brings in the relevant source text, making it much easier for models to match target sentences with their sources.
@@
Open-Source vs. Closed-Source. Among open-source models, parameter sizes vary widely, from a few billion (e.g., 6B–14B) to the much larger Deepseek V3 with 671B parameters. Despite these differences, larger models generally perform better in direct prompting and retrieval-augmented settings, especially in relationship classification. Deepseek-V3 (DP) shows strong performance with an F1 score of 47.59, outperforming many smaller models. However, when retrieval is applied, models like Qwen2.5-14B begin to reduce the gap with leading closed-source systems. For closed-source models, Gemini-1.5-Pro (RA) and GPT-4o (RA) achieve the highest F1 scores at 63.36 and 62.02, performing well in both source tracing and relationship classification. However, Gemini-1.5-Pro struggles with direct prompting, with an F1 score of only 12.29, highlighting the importance of retrieval. While closed-source models still lead overall, their advantage is reduced significantly when open-source LLMs use strong retrieval methods.
Figure 5: The performance of different models varies on different scenarios and source lengths.
@@ 5.4 Analysis
Figure 5 depicts the performance of different models across scenario and source length aspects.
Models Performance Across Scenarios. Under the DP method, GPT-4o performs better than all other models across all scenarios. With the RA method, Kimi, GPT-4o, and Gemini-1.5-Pro each exhibit distinct advantages in different scenarios. For instance, GPT-4o leads in EventSum, QA3doc, and QA2doc; Gemini-1.5-Pro outperforms others in MultiNews, MultiNews_e, QMSum, and LongSum; and Kimi shows outstanding performance in SAMSum and QA1doc. These results suggest that each model adapts differently depending on the scenario. They also show that RA leads to more significant improvements than DP, especially in multi-document, dialogue, and meeting scenarios.
Models Performance Across Source Lengths. Regarding source length, the RA is generally less affected by longer texts. In the DP, once the source length reaches 32k, only GPT-4o and DeepSeek maintain a passable but somewhat lower level of performance, while the others see a significant drop. Interestingly, Qwen25_14b usually falls behind Kimi and Gemini, but it surpasses both when the source length exceeds 20k.
@@ Appendix B Detail Experiment Results
In single-document English tasks, among open-source models, Qwen2.5-7B with retrieval-augmented tracing achieves the highest F1 (53.99), outperforming other open-source alternatives (e.g., Qwen2.5-14B with 49.19). However, the closed-source Kimi obtains an even higher F1 of 61.39 with retrieval, making it the top performer overall in this single-document English scenario. Notably, GPT-4o is quite capable under direct prompting (42.49), exceeding the direct-prompting baselines of most open-source LLMs. However, almost all models (open or closed) show significant gains when retrieval is introduced.
@@
In single-document Chinese tasks, among open-source models, DeepSeek-V3 (RA) leads with an F1 score of 60.43, outperforming Qwen and Llama. Among closed-source models, GPT-4o (RA) scores 55.54, while Gemini-1.5-Pro (RA) gets higher at 57.01. Although these closed-source models exceed most open-source options except DeepSeek-V3, GPT-4o also performs well without retrieval, scoring 53.21 with direct prompting. In contrast, Gemini relies heavily on retrieval, as shown by its sharp jump from a very low direct prompting score of 3.94 to 57.01 when retrieval is applied. This highlights the varying levels of dependence on reducing candidates among different models.
@@
In multi-document English tasks, Qwen1.5-14B (RA) leads among open-source models with an F1 score of 56.52, slightly ahead of DeepSeek-V3 (54.65). However, the closed-source Gemini-1.5-Pro gets the top score with 64.07, outperforming GPT-4o (60.24) and Kimi (62.87). GPT-4o also shows strong performance without retrieval, scoring 50.00, while Gemini struggles with a much lower 19.24. This suggests that GPT-4o is naturally better at direct prompting, whereas Gemini and Kimi depend more on retrieved-context to handle complex multi-document provenance.
@@
In multi-document Chinese tasks, DeepSeek-V3 (RA) leads open-source models with an F1 score of 72.93, far ahead of Qwen2.5-14B (49.39). However, GPT-4o achieves the best overall with 75.35, just ahead of Kimi (75.04) and Gemini-1.5-Pro (72.66). GPT-4o performs better in direct prompting than retrieval augmented tracing, scoring 77.72, much higher than Gemini and Kimi. This highlights its strong ability to handle multi-source Chinese text.
@@
In all, in single-document tasks, Qwen2.5-7B and DeepSeek-V3 emerge as strong open-source choices for English and Chinese, respectively, yet Kimi and Gemini can outperform them once retrieval is incorporated. GPT-4o stands out for its relatively high direct-prompting scores across both languages, showing strong built-in tracing capabilities. Under multi-document conditions, the complexity increases, and the top results often come from closed-source solutions (e.g., Gemini, GPT-4o, Kimi), although Qwen1.5-14B and DeepSeek-V3 hold their own in the open-source domain. Models integrating retrieval, whether open- or closed-source, generally exhibit greater gains and more accurate sentence-level provenance.
Appendix C Related Prompts

### v3

@@ Abstract
LLMs have achieved remarkable fluency and coherence in text generation, yet their widespread adoption has raised concerns about content reliability and accountability. In high-stakes domains, it is crucial to understand where and how the content is created. To address this, we introduce the Text pROVEnance (TROVE) challenge, designed to trace each sentence of a target text back to specific source sentences within potentially lengthy or multi-document inputs. Beyond identifying sources, TROVE annotates the fine-grained relationships (quotation, compression, inference, and others), providing a deep understanding of how each target sentence is formed. To benchmark TROVE, we construct our dataset by leveraging three public datasets covering 11 diverse scenarios (e.g., QA and summarization) in English and Chinese, spanning source texts of varying lengths (0-5k, 5-10k, 10k+), emphasizing the multi-document and long-document settings essential for provenance. To ensure high-quality data, we employ a three-stage annotation process: sentence retrieval, GPT-4o provenance, and human provenance. We evaluate 11 LLMs under direct prompting and retrieval-augmented paradigms, revealing that retrieval is essential for robust performance, larger models perform better in complex relationship classification, and closed-source models often lead, yet open-source models show significant promise, particularly with retrieval augmentation. We make our dataset available here: https://github.com/ZNLP/ZNLP-Dataset.
@@ 1 Introduction
Specifically, we first select examples with multi-sentence outputs to ensure sufficient sources and categorize data by original scenarios (question answering or summarization), languages (Chinese or English), input text length (0-5k, 5k-10k, or 10k+), and number of input documents (single-document or multi-document), aiming for a balanced distribution across each dimension. Next, we employ multiple information retrieval methods to recall the candidate source sentences for each sentence in the target text. Then, we use crafted prompts to guide GPT-4o in preliminary annotation, aiming to identify the sources of target sentences from retrieved candidate sentences and classify the target-source relationships into quotation (verbatim or partial copy), compression (summarization or paraphrase), inference (expansion, generalization, or specification), and others (e.g., negation). Each target sentence can be traced to multiple source sentences, and different relationships may apply simultaneously. For instance, a target sentence sourced from [a, b, c] might have [a, b] labeled as compression and [c] as quotation. Finally, we conduct a human review, requiring annotators to verify each annotation by considering both the source document(s) and the results produced by GPT-4o.
@@ 3 Task Formulation
Figure 2: Overview of our data annotation. (a) Sentence Retrieval: selecting candidate provenance sentences using multiple retrievers; (b) GPT-4o Provenance: automatically annotating provenance relationships based on retrieved sentences; (c) Human Provenance: reviewing and refining GPT-4o’s annotations while independently checking source documents to identify missing provenance sentences. Di-j denotes the j-th sentence in the i-th document.
@@ 4.2 Data Annotation
We employ GPT-4o to alleviate the manual annotation workload, as shown in Figure 2. For each target sentence $t_{i}$, the annotation procedure consists of three steps: (a) sentence retrieval, (b) GPT-4o provenance, and (c) human provenance.
Sentence Retrieval. Due to the lengthy input text, GPT-4o may ignore key sentences when directly tracing provenance through long passages. To mitigate this, we first retrieve candidate provenance sentences using the target sentences as queries and then perform provenance based on these sentences. We have presented ablation studies to validate this approach. Moreover, different retrievers capture diverse semantic features. To maximize the recall rate of candidate provenance sentences, we aggregate the results from $M$ distinct retrievers. Each retriever selects the top-$k$ most relevant sentences, denoted as $R_{i}(D,t_{i})$, forming the following set of candidate provenance sentences:
@@
GPT-4o Provenance. Based on the candidate provenance sentences, GPT-4o conducts fine-grained annotation and classifies the provenance relationship types, as depicted in Figure 2(b). The detailed prompt is provided in the appendix.
Human Provenance. Annotators review GPT-4o’s results to verify the provenance sentences and their corresponding relationship types. Noting that GPT-4o can ignore critical details, the annotators examine the document collections to address any omissions. As illustrated in Figure 2(c), the sentences “D2-2” and “D2-4”, which contain “12,000 books” and “digital reading machine” respectively, exhibit a strong connection to the target sentence, yet GPT-4o fails to identify them. Therefore, the annotators will incorporate these two sentences into the final analysis. We invite 8 graduate students to spend about 510 hours annotating the provenance of 4,388 English and 811 Chinese sentences, costing an average of \$0.20 per sentence.
@@ 4.4 Consistency Analysis
#Docs Lang Trace Type GPT-4o
@@
To ensure dataset quality, 10% of the examples are assigned to different annotators for consistency assessment. We evaluate annotator agreement from three perspectives: (1) tracing provenance sentences, (2) classification of relationship types, and (3) determination of necessary corrections to GPT-4o’s provenance sentences. To quantify agreement, we use Fleiss’ Kappa (Falotico and Quatto, 2015) to measure the reliability across multiple annotators. The results, presented in Table 2, demonstrate that the annotation process is reliable.
@@ 5.1 Experimental Setup
During GPT-4o provenance, initial retrieval significantly enhances GPT’s recall rate. Thus, each model is evaluated under two approaches: (1) direct prompting tracing, where the model processes the input directly, and (2) retrieval-augmented tracing, where retrieval is performed first, followed by tracing based on the retrieved results.
@@ 5.3 Evaluation Results
Vicuna-7b DP 06.80 23.22 10.50 02.37 08.53 3.69 06.78 23.10 10.44 02.38 08.43 3.69 7.08
RA 27.14 41.76 32.50 9.79 17.22 12.38 29.10 40.78 33.64 10.39 15.76 12.44 22.74
LLama3-8b DP 05.16 16.63 06.97 02.03 07.21 02.96 05.45 15.50 06.43 01.84 06.20 02.49 04.71
RA 43.74 38.19 40.61 22.49 19.51 20.82 49.81 35.04 41.07 25.42 18.40 21.33 30.96
Chatglm-6b DP 00.02 00.04 00.02 00.01 00.00 00.01 00.04 00.02 00.03 00.01 00.01 00.01 0.02
RA 03.68 04.06 03.84 01.50 01.76 01.60 11.97 03.98 05.93 04.85 01.73 02.53 3.47
Qwen1.5-7b DP 06.47 41.00 11.18 01.36 11.71 02.36 06.50 40.82 10.84 01.27 10.04 02.20 6.65
RA 35.99 53.25 42.26 11.02 19.08 13.80 34.83 52.19 41.00 10.26 16.48 12.48 27.39
Qwen2.5-7b DP 08.88 49.56 14.77 02.89 15.49 4.81 07.67 49.50 12.94 01.93 13.56 3.29 8.95
RA 41.96 71.92 52.23 14.73 28.56 19.23 39.50 69.72 49.65 12.32 24.55 16.21 34.33
Qwen1.5-14b DP 12.50 38.97 16.84 03.71 13.35 05.25 13.36 37.51 16.93 03.65 11.32 04.74 10.94
RA 45.07 53.68 48.33 14.06 20.96 16.72 47.99 51.65 49.20 14.50 17.90 15.93 32.54
Qwen2.5-14b DP 29.77 52.25 36.12 15.11 27.37 18.69 29.50 49.55 33.01 14.37 24.74 16.24 26.02
RA 54.60 50.24 51.99 29.22 27.43 28.12 64.68 47.02 54.23 33.49 24.80 28.37 40.68
DeepSeek-V3 (671B) DP 44.79 56.56 49.80 21.88 28.40 24.63 39.54 54.31 44.95 17.46 25.61 20.41 34.95
RA 49.17 55.85 51.94 26.10 31.20 28.24 50.75 53.76 51.93 26.17 28.54 27.19 39.83
@@
Impact of Retrieval-Augmented Tracing vs. Direct Prompting Tracing. Across almost all models, retrieval-augmented tracing outperforms direct prompting in F1 scores, often by a large margin. For example, Qwen2.5-14B’s F1 jumps from 26.02 to 40.68 with retrieval, while ChatGLM-6B, which nearly fails in direct prompting tracing with an F1 of 0.02, improves to 3.47. Even closed-source models show the same trend, as Gemini-1.5-Pro significantly improves from an F1 of 9.61 to 51.18 with retrieval. It suggests that retrieval helps overcome context-length limits and brings in the relevant source text, making it much easier for models to match target sentences with their sources.
@@
Open-Source vs. Closed-Source. Among open-source models, parameter sizes vary widely, from a few billion (e.g., 6B–14B) to the much larger Deepseek V3 with 671B parameters. Despite these differences, larger models generally perform better in direct prompting and retrieval-augmented settings, especially in relationship classification. Deepseek-V3 (DP) shows strong performance with an F1 score of 34.95, outperforming many smaller models. However, when retrieval is applied, models like Qwen2.5-14B begin to reduce the gap with leading closed-source systems. For closed-source models, Gemini-1.5-Pro (RA) and GPT-4o (RA) achieve the highest F1 scores at 51.18 and 48.97, performing well in both source tracing and relationship classification. However, Gemini-1.5-Pro struggles with direct prompting, with an F1 score of only 9.61, highlighting the importance of retrieval. While closed-source models still lead overall, their advantage is reduced significantly when open-source LLMs use strong retrieval methods.
Figure 5: The performance of different models varies on different scenarios.
Figure 6: The performance of different models varies on different sourth lengths.
@@
(a) Confusion matrix for GPT-4o (DP).
(b) Confusion matrix for GPT-4o (RA).
Figure 7: Confusion matrix for GPT-4o. X-axis: predictions, y-axis: ground truth.
@@ 5.4 Analysis
Figure 5 and Figure 6 show model performance across scenarios and source lengths respectively.
Models Performance Across Scenarios. Under the DP method, GPT-4o performs better than all other models across all scenarios. With the RA method, Kimi, GPT-4o, and Gemini-1.5-Pro each exhibit distinct advantages in different scenarios. For instance, GPT-4o leads in EventSum, QA3doc, and QA2doc; Gemini-1.5-Pro outperforms others in MultiNews, MultiNews_e, QMSum, QA1doc, and LongSum; and Kimi shows outstanding performance in SAMSum. These results suggest that each model adapts differently depending on the scenario. They also show that RA leads to more significant improvements than DP, especially in multi-document, dialogue, and meeting scenarios.
Models Performance Across Source Lengths. Regarding source length, the RA is generally less affected by longer texts. In the DP, once the source length reaches 32k, only GPT-4o and DeepSeek-V3 maintain a passable but somewhat lower level of performance, while the others see a significant drop. Interestingly, Qwen2.5-14b usually falls behind Kimi and Gemini-1.5-Pro, but it surpasses both when the source length exceeds 20k.
Error Analysis. To understand model behavior, we analyze the confusion matrices for GPT-4o under both direct prompting and retrieval-augmented conditions, with results shown in 7(a) and 7(b) respectively. The matrices reveal a clear hierarchy of relationship difficulty: Quotation is easiest, followed by Compression, while Inference proves most challenging. Three error patterns emerge: (1) Compression bias — models overpredict this category, with 533 Quotation and 311 Inference instances misclassified as Compression in RA; (2) "Inference→Compression" confusion — 311 out of 672 true Inference cases are misclassified as Compression, indicating difficulty distinguishing between summarization and logical derivation; (3) "Other" underrepresentation — only 7 out of 28 instances correctly identified, highlighting challenges with rare relationship types.
Impact of Retrieval. While RA improves overall performance (Inference correct predictions increase from 168 to 299), it also intensifies misclassification attempts. The Quotation→Compression errors increase from 401 in DP to 533 in RA, suggesting that additional context sometimes causes models to overinterpret simple quotations as more complex relationships. These patterns reveal fundamental challenges in relationship classification that extend beyond performance metrics.
@@ 6 Conclusion
Acknowledgements
The research work described in this paper has been partially supported by the Natural Science Foundation of China under Grant No. 62336008. We sincerely thank the anonymous reviewers for their insightful comments and constructive suggestions.
@@ Appendix B Detail Experiment Results
In single-document English tasks, among open-source models, Qwen2.5-7B with retrieval-augmented tracing achieves the highest F1 (35.02), outperforming other open-source alternatives (e.g., Qwen2.5-14B with 34.50). However, the closed-source Gemini-1.5-Pro obtains an even higher F1 of 48.39 with retrieval, making it the top performer overall in this single-document English scenario. Notably, GPT-4o is quite capable under direct prompting (33.97), exceeding the retrieval-augmented baselines of most open-source LLMs. However, almost all models (open or closed) show significant gains when retrieval is introduced.
@@
In single-document Chinese tasks, among open-source models, DeepSeek-V3 (RA) leads with an F1 score of 44.52, outperforming Qwen and Llama. Among closed-source models, GPT-4o (DP) scores 42.99, while Gemini-1.5-Pro (RA) gets higher at 46.95. Although these closed-source models exceed most open-source options except DeepSeek-V3, GPT-4o also performs well without retrieval, scoring 43.61 with direct prompting, even better than its RA variant. In contrast, Gemini relies heavily on retrieval, as shown by its sharp jump from a very low direct prompting score of 2.85 to 46.95 when retrieval is applied. This highlights the varying levels of dependence on reducing candidates among different models.
@@
In multi-document English tasks, Qwen2.5-14b (RA) leads among open-source models with an F1 score of 44.54, slightly ahead of DeepSeek-V3 (43.39). However, the closed-source Gemini-1.5-Pro gets the top score with 51.25, outperforming GPT-4o (48.34) and Kimi (49.48). GPT-4o also shows strong performance without retrieval, scoring 41.26, while Gemini struggles with a much lower 15.80. This suggests that GPT-4o is naturally better at direct prompting, whereas Gemini and Kimi depend more on retrieved context to handle complex multi-document provenance.
@@
In multi-document Chinese tasks, DeepSeek-V3 (RA) leads open-source models with an F1 score of 54.52, far ahead of Qwen2.5-14B (47.51). However, GPT-4o achieves the best overall with 61.09, just ahead of Gemini-1.5-Pro (58.11) and Kimi (57.33). This highlights GPT-4o’s strong ability to handle multi-source Chinese text.
@@
In all, in single-document tasks, Qwen2.5-7B and DeepSeek-V3 emerge as strong open-source choices for English and Chinese, respectively, yet Gemini-1.5-Pro can outperform them once retrieval is incorporated. GPT-4o stands out for its relatively high direct-prompting scores across both languages, showing strong built-in tracing capabilities. Under multi-document conditions, the complexity increases, and the top results often come from closed-source solutions (e.g., Gemini-1.5-pro, GPT-4o, Kimi), although Qwen2.5-14b and DeepSeek-V3 hold their own in the open-source domain. Models integrating retrieval, whether open- or closed-source, generally exhibit greater gains and more accurate sentence-level provenance.
B.1 Confusion Matrix Analysis Details
The observed discrepancies in total counts between confusion matrices for different methods are attributed to the following methodological factors: (1) The retrieval-augmented method may fail to retrieve certain sentences from the source documents, leading to variations in the number of ground-truth sentences available for classification. (2) The retrieval-augmented and direct prompting methods trace different sets of source sentences due to their distinct retrieval mechanisms. Sentences that remain untraced by either method are excluded from the subsequent relationship classification task, resulting in different sample sizes across experimental conditions. It is worth noting that we employ the pass@5 evaluation metric for all experimental assessments to ensure consistent and robust performance measurement.
Appendix C Prompts